# The concept of minimal dissipation and the identification of work in autonomous systems: A view from classical statistical physics

### v1

@@ I Introduction
Work is a central concept both in thermodynamics and in mechanics. While the definition of the mechanical work done on an isolated, classical many-body system is straightforward, the notion of thermodynamic work is more involved. In thermodynamics, the definition of work requires a distinction between a system and its environment. Further, the properties of the environment need to be specified as well as conditions on the strength of the coupling to the system [1, 2, 3, 4]. If the coupling is strong, the system and the environment are correlated or the environment contains only few degrees of freedom, it is unclear how to define work and whether it is a meaningful concept at all.  [^1] This statement holds for classical as well as quantum mechanical systems.
One requirement a definition of work needs to fulfil is that in the limit of a quasi-static process performed on a system coupled to a heat bath it reproduces the definition given in equilibrium thermodynamics. (In this article we use the term heat bath for an environment with infinite heat capacity, which is weakly coupled to the system of interest, while we use the term environment for any type of system coupled in any way to the system of interest.) I.e. if work is done on a system by externally changing some of its parameters with time, such as the strength of a magnetic field or the volume available to the system, and if these parameters are varied infinitely slowly and the system is in contact with a heat bath, the total work is the difference between the final and the initial equilibrium free energy of the system: $W=\Delta F=F_{f}-F_{i}$. [6, 4] In such a quasi-static transformation the system remains in equilibrium with the bath and the total entropy does not change [1]. If the parameters are changed externally at a non-zero rate, the process is in general irreversible and $W$ will, on average, exceed the free energy difference $\langle W\rangle\geq\Delta F$. [6] Or conversely, if we wish to extract work from the system, it will be less than the available free energy.
@@
We will use the framework of autonomous systems. I.e. instead of imposing an external driving force on the Hamiltonian or coupling the system to a specific type of bath, we consider a composite, isolated supersystem made of the system $S$ and its environment $E$. This composite supersystem is governed by a Hamiltonian that is independent of time. We obtain the dynamics of $S$ by tracing out the degrees of freedom of $E$. The resulting equation of motion (EoM) for the density of states of the system can in general be written as a term containing a commutator with an effective, time-dependent Hamiltonian, which in the literature is often called the conservative part, and a rest called the dissipative part [14, 15, 16] (we will see later that these names can be misleading, as the interpretation in terms of dissipation is not always given). To systematically integrate out degrees of freedom of an autonomous system has the advantage over other approaches, that the setting is as general as possible, i.e. a priori neither assumptions on the coupling between the system and the environment nor on the properties of the environment are required. Several authors have brought forward definitions of work based on the effective Hamiltonian in such a setting [17, 18, 19, 20, 14, 21]. However, as the splitting between the conservative and the dissipative part is not unique, different authors have suggested different identifications of internal energy, work and heat [13]. Here we will analyse which type of splitting produces an effective Hamiltonian that equals the free energy in the limit of a quasi-static process.
@@
We will begin our discussion with classical supersystems. If one uses the projection operator introduced by Zwanzig [22] to integrate out the degrees of freedom of the environment then, in the classical case, one obtains an EoM for the system which contains the free energy in the conservative part, i.e. in a Poisson bracket [15]. We briefly recall Izvekov’s derivation of this equation. Then we show that the EoM for the corresponding relevant density has a similar structure. We argue that this EoM is equal to the one obtained via the principle of minimal dissipation, however this requires a different choice of inner product than used in the original work by Colla and Breuer [14]. Finally, we show that this line of reasoning cannot be extended to open quantum system in a general manner, but that it does hold for systems and environments with equilibrium states that factorize.
@@ II Classical-quantum analogy
This expression is formulated in the Schrödinger picture, i.e., the time-dependence is carried by the density. Equivalently the Heisenberg picture can be used, where the time-dependence is carried by the observables and the average is taken with respect to the initial density $\langle B\rangle_{t}=\text{Tr}(\rho(0)B_{H}(t))$. Since these expectation values are the same, the Liouville equation for the observable $B_{H}$ is
@@
[…]
@@ II.1 Hamiltonian of mean force
The Hamiltonian of mean force is a central concept in the formulation of fluctuation relations [23, 24]. We briefly recall its definition: The Hamiltonian of a composite system can be decomposed into components that act on the subsystem of interest only, the environment only, and on the interaction between them, i.e. $H=H_{S}+H_{E}+H_{SE}$. In thermal equilibrium the global state is represented by a state $\varrho_{\beta}=e^{-\beta H}/Z$ with the partition function $Z=\text{Tr}(e^{-\beta H})$. The reduced equilibrium state of the subsystem is then found by tracing over the environmental degrees of freedom
@@
The Hamiltonian of mean force $H_{S}^{\ast}$ is introduced as the effective Hamiltonian that describes the equilibrium of the reduced system. (The term Hamiltonian of mean force is used predominantly in the context of quantum mechanics, while the same quantity is called free energy landscape or effective free energy in physical chemistry. In the context of molecular modelling of biomolecules and polymers the related potential of mean force is often used, which is obtained if one traces not only over the environment but also over the momenta of the system particles.) Its classical and quantum form can be expressed in complete analogy using the the partial trace in its classical/quantum mechanical form
$H^{\ast}_{S}=-k_{B}T\ln\text{Tr}_{E}(e^{-\beta H})/\text{Tr}_{E}(e^{-\beta H_{E}})\;.$ (10)
Note, eq. (9) only determines $H_{S}^{\ast}$ up to an additive constant and the expression (10) is based on the common choice $Z_{S}^{\ast}=Z/Z_{E}$ [25]. The free energy of $S$ is then obtained via the partition function
$F_{S}=-k_{B}T\ln Z_{S}^{\ast}=-k_{B}T\ln\text{Tr}_{S}(e^{-\beta H_{S}^{\ast}})$
@@ III Projection operator technique
Projection operator techniques can now be applied to derive EoM for macroscopic variables or for the degrees of freedom of the subsystem of interest by decomposing the dynamics. The first step is the definition of a projector that maps the space of relevant variables on itself. Such an operator can be defined by the choice of an appropriate scalar product or relevant density [26].
@@
Alternatively we can construct an EoM for the so called relevant density, i.e. the probability density associated with the relevant observables. Then we work with the adjoint projector defined through
@@ IV Zwanzig projector in classical statistical mechanics
The functions $\psi_{\alpha}$ fix the values of the relevant observables to numbers $\alpha_{k}$
$\psi_{\alpha}(\Gamma)=\delta(A(\Gamma)-\alpha)=\prod_{k}\delta(A_{k}(\Gamma)-\alpha_{k})\;.$ (24)
They form a set of functions with the continuous index $\alpha$ and they have the useful property
@@
Thus the mean value $p(\alpha,t)=\text{Tr}(\psi_{\alpha}\rho(t))$ defines the ”macroscopic probability density” of the observables $A$, i.e. $p(\alpha,t)\text{d}\alpha$ is the probability to find the values of the observables $A$ in the volume element $\text{d}\alpha$ around $\alpha$ if the ensemble is distributed according to $\varrho$.
The trace in the numerator of the second term of eq. (23) integrates $B(\Gamma)$ over all microstates $\Gamma$ for which $A(\Gamma)=\alpha$, where the microstates are weighted according to the probability distribution $\varrho(\Gamma)$. Hence the Zwanzig projector contains a conditional probability in the ensemble specified by $\varrho(\Gamma)$. It maps the observable $B$ to the best possible approximation of $B$ in terms of functions of $A$ [29]. This is the property we need in order to define a Hamiltonian of mean force.
The corresponding adjoint projector defined by eq. (15) is given by
@@ IV.1 Dirft Term, Conservative Force
Depending on the context, in the literature on classical systems this term is sometimes called ”drift” [27, 24] and sometimes ”conservative force” [30]. We now set the weight in the projector to the canonical equilibrium distribution $\varrho=\varrho_{\beta}(\Gamma)=e^{-\beta H(\Gamma)}/\text{Tr}(e^{-\beta H})$. Then we can exploit the fact that for any observable, and particularly $A$:
@@
If $A$ is a part of a set of canonical variables (e.g., all positions and momenta of the particles in $S$) we can use $\{A,\,\cdot{}\}_{\Gamma}=\{A,\,\cdot\,\}_{A}=-\mathbb{J}\nabla_{A}\,\cdot\,$.
The numerator on the right of eq. (30) becomes $-k_{B}T\text{Tr}(\psi_{\alpha}\{A,\rho_{\beta}\})$. The propagator only acts on the $\psi_{\alpha}(\Gamma)$ and yields $\exp(i\mathcal{L}t)f(A)=f(A_{H}(t))$. Thus,
@@
$=$ $-k_{B}T\mathbb{J}\nabla_{A}\ln\text{Tr}(\varrho_{\beta}\psi_{A_{H}(t)})$ (35)
=:absent:=:= : $\{A,H^{\ast}(A,t)\}.$ (36)
In the last step we have identified the effective free energy associated with the values of $A$ taken in the equilibrium ensemble $H^{\ast}(A,t)=-k_{B}T\ln\text{Tr}(\varrho_{\beta}\psi_{A_{H}(t)})$. If the system undergoes a quasistatic process from a macrostate with one set of values of $A$ to another, the difference in the effective free energy equals the work performed on the system. So by choosing the appropriate projection operator it is indeed possible to construct an EoM, in which the equilibrium work appears in the Poisson bracket. In the literature on physical chemistry, the term $-k_{B}T\ln\text{Tr}(\varrho_{\beta}\psi_{A_{H}(t)})$ is usually denoted by $\Delta F(A)$ or $\Delta G(A)$ to emphasize the relation to a thermodynamic potential. Here we used a different notation, because this term turns into a Hamiltonian of mean force if we specify $A$ to be the canonical degrees of freedom of the system.
@@
In summary, so far we have integrated out the degrees of freedom of the environment without making any approximations, and we have obtained an equation of motion for the system, which contains a Poisson bracket with the Hamiltonian of mean force, i.e. with the quantity that is equal to work in the quasi-static case [15].
@@
With $\psi_{\alpha}\psi_{\alpha^{\prime}}=\psi_{\alpha}\delta(\alpha-\alpha^{\prime})$ we get
@@
In both cases this fraction was turned into the effective free energy in the following steps. Accordingly, the drift term in the EoM for $p(\alpha,t)$ also features a Poisson bracket structure with $H^{\ast}(\alpha)$. (Note, it is crucial that $A$ represents a canonical set to obtain the symplectic structure.)
@@ V Relation to Minimal Dissipation
$\dot{\rho}_{S}(t)=-i[K_{S}(t),\rho_{S}(t)]+\mathcal{D}_{t}[\rho_{S}(t)]\;,$ (49)
are often used, where $K_{S}$ is the effective Hamiltonian, $\mathcal{D}_{t}$ is a dissipator of the generalized Lindblad form and $\rho_{S}$ is the density matrix of the system. The function $p(\alpha,t)$ defined in sec. IV is the classical equivalent of $\rho_{S}$ if the observables $A$, which are set to the values $\alpha$ in $p(\alpha,t)$, are the canonical degrees of freedom of the system [31]. In the form of eq. (49) the decomposition of the master equation in a conservative and a dissipative part is not unique. A distinct splitting can be achieved by specifying a norm on the space of dissipative superoperators and defining the effective Hamiltonian to be the one whose corresponding dissipator is minimal [32]. This Hamiltonian is used to define work in ref. [14, 21, 33].
@@
The condition on minimal dissipation is not the source of the discrepancy, either. We did not impose the condition explicitly, however, our derivation fulfils it by construction. The splitting between the conservative term and the dissipative term in eq. (49) as well as in eq. (22) is determined by the projection operator. Depending on the choice of the functions $\{\phi_{k}\}$ in eq. (11) contributions to the dynamics get shuffled from one part of eq. (12) to the other. One extreme case would be the Mori-projection operator [36] which projects onto only one observable, i.e. the sum in eq. (11) runs over only one function $\phi_{1}$. The Zwanzig projection operator is the opposite extreme case, because it requires a complete basis. It is this requirement that implicitly imposes the condition of minimal dissipation. To see this, we note that the term minimal here refers to the norm induced by the inner product. Under this norm,
@@
where $\mathcal{P}$ is any projection operator, $B$ is any observable and $\mathcal{P}^{Z}$ is the Zwanzig projection operator as defined in eq. (23). Thus the projection operator we chose in sec. IV.1 maximizes the drift term (i.e. the conservative term) and minimizes the rest (i.e. the dissipative term).
The crucial difference between sec. IV and ref. [14] is the definition of the inner product that induces the norm. The inner product employed in ref. [32] does not contain a weight, while the inner product required for the Hamiltonian of mean force to appear in eq. (44) contains the equilibrium measure $\varrho_{\beta}$. As demonstrated in ref. [37] a unique decomposition of a given generator can be achieved with respect to weighted scalar products. We thus agree with Colla and Breuer, one can use the generalized master equation and its unique splitting to define work, however, we propose to use a different inner product in order to obtain the correct limit for quasi-static processes performed on systems coupled to a heat bath.

### v2

@@ I Introduction
Work is a central concept both in thermodynamics and in mechanics. While the definition of the mechanical work done on an isolated, classical many-body system is straightforward, the notion of thermodynamic work is more involved. In thermodynamics, the definition of work requires a distinction between a system and its environment. Further, the properties of the environment need to be specified as well as conditions on the strength of the coupling to the system [1, 2, 3, 4]. If the coupling is strong, the system and the environment are correlated, or the environment contains only few degrees of freedom, it is unclear how to define work and whether it is a meaningful concept at all.  [^1] This statement holds for classical as well as quantum mechanical systems.
One requirement a definition of work needs to fulfil is that in the limit of a quasi-static process performed on a system coupled to a heat bath it reproduces the definition given in equilibrium thermodynamics. (In this article we use the term heat bath for an environment with infinite heat capacity, which is weakly coupled to the system of interest, while we use the term environment for any type of system coupled in any way to the system of interest.) I.e. if work is done on a system by externally changing some of its parameters (such as the strength of a magnetic field or the volume available to the system) and if these parameters are varied infinitely slowly and the system is in contact with a heat bath, the total work is the difference between the final and the initial equilibrium free energy of the system: $W=\Delta F=F_{f}-F_{i}$. [6, 4] In such a quasi-static transformation the system remains in equilibrium with the bath and the total entropy does not change [1]. If the parameters are changed faster, the process is in general irreversible and $W$ will, on average, exceed the free energy difference $\langle W\rangle\geq\Delta F$. [6] Or conversely, if we wish to extract work from the system, it will be less than the available free energy.
@@
We will use the framework of autonomous systems. In this framework, instead of imposing an external driving force on the Hamiltonian or coupling the system to a specific type of bath, one considers a composite, isolated supersystem made of the system $S$ and its environment $E$. This composite is governed by a Hamiltonian that is independent of time. In such a setting, the environment $E$ acts as a source of work for the system $S$, i.e. the time-dependent external driving, which is required for work to be performed on $S$, is produced by the dynamics of $E$. This is a standard approach used for example in ref. [15, 16, 17, 18, 19, 20, 14, 21, 22, 23].
We obtain the dynamics of $S$ by tracing out the degrees of freedom of $E$. The resulting equation of motion (EoM) for the density operator of $S$ can in general be written in the form
$\dot{\rho}_{S}(t)=-i[H^{\rm eff}_{S}(t),\rho_{S}(t)]+\mathcal{D}_{t}[\rho_{S}(t)]\;,$ (1)
i.e. in terms of a commutator that contains an effective, time-dependent Hamiltonian and a rest. In the case of a classical system, the density operator is replaced by the phase space probability density and the commutator by a Poisson bracket, but the structure of the equation remains the same. In the literature the first term is often called the conservative part and the rest the dissipative part [24, 14, 25, 26] (we will see later that these names can be misleading, as the interpretation in terms of dissipation is not always given). Work is then defined via the equation for the work flux related to the effective Hamiltonian
$\dot{W}_{S}(t)=\text{Tr}\Big{(}\dot{H}_{S}^{\rm eff}\,\rho_{S}\Big{)}\;.$ (2)
@@
To systematically integrate out degrees of freedom of an autonomous system has the advantage over other approaches, that the setting is as general as possible, i.e. a priori neither assumptions on the coupling between the system and the environment nor on the properties of the environment are required. Several authors have brought forward definitions of work based on the effective Hamiltonian in such a setting [16, 23, 21, 22, 14, 27]. However, as the splitting between the conservative part and the dissipative part is not unique, different authors have suggested different identifications of internal energy, work and heat [13]. Here we will analyse which type of splitting produces an effective Hamiltonian that equals the Hamiltonian of mean force in the limit of a quasi-static process.
@@
We will begin with an introduction to our notation that highlights the similarity between the classical and the quantum mechanical description of the dynamics. We proceed with the discussion of projection operator techniques applied to classical supersystems. If one uses the projection operator introduced by Zwanzig [28] to integrate out the degrees of freedom of the environment then, in the classical case, one obtains an EoM for the system which contains the Hamiltonian of mean force in the conservative part, i.e. in a Poisson bracket [25]. We briefly recall Izvekov’s derivation of this equation. Then we show that the EoM for the corresponding relevant density has a similar structure. We argue that this EoM is equal to the one obtained via the principle of minimal dissipation, however this requires a different choice of inner product than used in the original work by Colla and Breuer [14]. Finally, we show that this line of reasoning cannot be extended to open quantum systems in a general manner but that it does hold for systems and environments with equilibrium states that factorize.
@@ II Classical-quantum analogy
This expression is formulated in the Schrödinger picture, i.e., the time-dependence is carried by the density. Equivalently the Heisenberg picture can be used, where the time-dependence is carried by the observables and the average is taken with respect to the initial density $\langle B\rangle_{t}=\text{Tr}(\rho(0)B_{H}(t))$. Since these expectation values are the same, the Liouville equation for the observable $B_{H}(t)$ is
@@
$\text{ with }i\mathcal{L}_{H}X=\begin{cases}-\{H_{H},X\}\\ i[H_{H},X]\end{cases}.$ (10)
@@ II.1 Hamiltonian of mean force
We are interested in the description of a subsystem of a composite system, i.e., only the degrees of freedom of the subsystem are of interest and the rest is treated as the environment. The Hamiltonian of mean force is a useful analytical tool to treat the average effect of environmental degrees of freedom on variables of interest [29]. We briefly recall its definition: The Hamiltonian of a bipartite system has the general form $H=H_{S}+H_{E}+H_{SE}$, where $H_{S}$ acts only on the subsystem $S$ and $H_{E}$ acts only on the environment $E$. The third term $H_{SE}$ represents the interaction operator. A thermal equilibrium state at some inverse temperature $\beta$ is given by $\varrho_{\beta}=e^{-\beta H}/Z$ with the partition function $Z=\text{Tr}(e^{-\beta H})$. (Note that we do not imply here that the composite system was coupled to some additional external heat bath, we simply consider an isolated supersystem prepared in a state of the canonical form.) Then the reduced state of the system $S$ is found by tracing over the environmental degrees of freedom
@@
with the Hamiltonian of mean force
$H^{\ast}_{S}:=-k_{B}T\ln\text{Tr}_{E}(e^{-\beta H})/\text{Tr}_{E}(e^{-\beta H_{E}})\;.$ (12)
In general, $H^{\ast}_{S}$ differs from the bare system Hamiltonian $H_{S}$ [30]. Eq. (11) only determines $H_{S}^{\ast}$ up to an additive constant. Any addition would also change the partition function and leave averages unaffected. The expression (12) is based on the common choice $Z^{\ast}_{S}=Z/Z_{E}$ [30, 31].
The term Hamiltonian of mean force is used predominantly in the context of quantum mechanics. In physical chemistry, the same quantity is called effective free energy or restricted free energy [32, 25]. In the context of molecular modelling of biomolecules and polymers a related quantity is often used, the potential of mean force. The potential of mean force is obtained by tracing not only over the environment $E$ in eq. (12) but also over the momenta of the particles in $S$, i.e. over the contribution of the kinetic energy of the system.
The relation between the Hamiltonian of mean force and the thermodynamic work is made as follows: The free energy of $S$ is obtained by taking the logarithm of the partition function associated with the Hamiltonian of mean force
$F_{S}=-k_{B}T\ln Z_{S}^{\ast}=-k_{B}T\ln\text{Tr}_{S}(e^{-\beta H_{S}^{\ast}})\;,$ (13)
i.e. the Hamiltonian of mean force is directly related to the free energy. If the energy contained in the coupling term $H_{SE}$ is small compared to $H_{S}$ and if the environment is sufficiently large to stay in a canonical state, then the free energy difference equals the work performed in a quasi-static process.
@@ III Projection operator technique
Projection operator techniques can be applied to derive EoM for macroscopic variables or for the degrees of freedom of the subsystem of interest by decomposing the dynamics. The first step is the definition of a projector that maps the space of relevant variables on itself. Such an operator can be defined by the choice of an appropriate scalar product or a relevant density [33].
@@
Alternatively, we can construct an EoM for the so-called relevant density, i.e. the probability density associated with the relevant observables. Then we work with the adjoint projector defined through
@@ IV Zwanzig projector in classical statistical mechanics
To define the Zwanzig projector, we first introduce functions $\psi_{\alpha}$ which fix the values of the relevant observables $A_{k}$ to the numerical values $\alpha_{k}$.
$\psi_{\alpha}(\Gamma):=\delta(A(\Gamma)-\alpha)=\prod_{k}\delta(A_{k}(\Gamma)-\alpha_{k})\;.$ (26)
These form a set of functions with the continuous index $\alpha$ and with the useful property
@@
The value $p(\alpha,t)=\text{Tr}(\rho(t)\psi_{\alpha})$ defines the so-called ”macroscopic probability density” of the observables $A$, i.e. in an ensemble that is distributed according to $\varrho$, $p(\alpha,t)\text{d}\alpha$ is the probability to find the values of the observables $A$ in the volume element $\text{d}\alpha$ around the values $\alpha$.
The trace in the numerator of the second term of eq. (28) integrates $B(\Gamma)$ over all microstates $\Gamma$ for which $A(\Gamma)=\alpha$, where the microstates are weighted according to the probability distribution $\varrho(\Gamma)$. Hence, the Zwanzig projector contains a conditional probability in the ensemble specified by $\varrho(\Gamma)$. It maps the observable $B$ to the best possible approximation of $B$ in terms of functions of $A$ [36]. This is the property we need in order to define a Hamiltonian of mean force.
@@
The adjoint projector defined by eq. (18) is given by
@@ IV.1 Dirft Term, Conservative Force
Depending on the context, in the literature on classical systems this term is sometimes called ”drift” [34, 32] and sometimes ”conservative force” [37]. We now set the weight in the projector to the canonical equilibrium distribution $\varrho=\varrho_{\beta}(\Gamma)=e^{-\beta H(\Gamma)}/\text{Tr}(e^{-\beta H})$. (Note that this does not imply that the system were coupled to a heat bath. We simply choose a weight to define a specific projection operator, but we still consider the equations of motion for the general case.) Then we can exploit the fact that for any observable, and particularly $A$:
@@
If $A$ is a subset of a set of canonical variables (e.g., all positions and momenta of the particles in $S$) we can use $\{A,\,\cdot{}\}_{\Gamma}=\{A,\,\cdot\,\}_{A}=-\mathbb{J}\nabla_{A}\,\cdot\,$.
The numerator on the right of eq. (33) becomes $k_{B}T\text{Tr}(\psi_{\alpha}\{A,\rho_{\beta}\})$. The propagator only acts on the $\psi_{\alpha}(\Gamma)$ and yields $\exp(i\mathcal{L}t)f(A)=f(A_{H}(t))$. Thus,
@@
$=$ $-k_{B}t\mathbb{J}\nabla_{A}\ln\int\text{d}\Gamma\varrho_{\beta}(\Gamma)\delta(A(\Gamma)-A(\Gamma(t))$ (38)
$=:$ $-k_{B}T\mathbb{J}\nabla_{A}\ln\text{Tr}(\varrho_{\beta}\psi_{A_{H}(t)})$ (39)
$=:$ $\{A,H^{\ast}(A,t)\}.$ (40)
In the last step, we have identified the Hamiltonian of mean force associated with the values of $A$ taken in the equilibrium ensemble $H^{\ast}(A,t)=-k_{B}T\ln\text{Tr}(\varrho_{\beta}\psi_{A_{H}(t)})$. As shown in sec. II.1, the free energy is obtained by tracing over $H^{\ast}(A,t)$. So by choosing the appropriate projection operator it is indeed possible to construct an EoM, in which the effective Hamiltonian that appears in the Poisson bracket has a direct connection to the thermodynamic potential and can be used to determine work done by the mean forces. If the environment remains in a canonical state and evolves slowly, then both the temperature and the mean force applied to the subsystem would change quasistatically.
@@
In summary, so far we have integrated out the degrees of freedom of the environment without making any approximations, and we have obtained an equation of motion for the system, which contains a Poisson bracket with the Hamiltonian of mean force, i.e. with the quantity that is directly related to the thermodynamic equilibrium work [25].
@@
With $\psi_{\alpha}\psi_{\alpha^{\prime}}=\delta(\alpha-\alpha^{\prime})\psi_{\alpha}$ we get
@@
In both cases this fraction was turned into the Hamiltonian of mean force in the following steps. Accordingly, the drift term in the EoM for $p(\alpha,t)$ also features a Poisson bracket structure with $H^{\ast}(\alpha)$. (Note, it is crucial that $A$ represents a canonical set to obtain the symplectic structure.)
@@ V Relation to Minimal Dissipation
$\dot{\rho}_{S}(t)=-i[H^{\rm eff}_{S}(t),\rho_{S}(t)]+\mathcal{D}_{t}[\rho_{S}(t)]\;,$ (53)
are often used, where $H^{\rm eff}_{S}$ is the effective Hamiltonian, $\mathcal{D}_{t}$ is a dissipator of the generalized Lindblad form and $\rho_{S}$ is the density matrix of the system. The function $p(\alpha,t)$ defined in sec. IV is the classical equivalent of $\rho_{S}$ if the observables $A$, which are set to the values $\alpha$ in $p(\alpha,t)$, are the canonical degrees of freedom of the system [38]. In the form of eq. (53) the decomposition of the master equation in a conservative and a dissipative part is not unique. A distinct splitting can be achieved by generalizing the Hilbert-Schmidt scalar product for operators to superoperators and minimizing the dissipator with respect to the induced norm [39]. The resulting effective Hamiltonian is used to define work in ref. [14, 27, 40].
@@
The condition of minimal dissipation is not the source of the discrepancy, either. We did not impose the condition explicitly, however, our derivation fulfils it by construction. The splitting between the conservative term and the dissipative term in eq. (53) as well as in eq. (25) is determined by the projection operator. Depending on the choice of the functions $\{\phi_{k}\}$ in eq. (14) contributions to the dynamics get shuffled from one part of eq. (15) to the other. One extreme case would be the Mori projection operator [43] which projects onto only one observable, i.e. the sum in eq. (14) runs over only one function $\phi_{1}$. The Zwanzig projection operator is the opposite extreme case, because it requires a complete basis. It is this requirement that implicitly imposes the condition of minimal dissipation. To see this, we note that the term minimal here refers to the norm induced by the inner product. Under this norm,
@@
where $\mathcal{P}$ is any projection operator, $B$ is any observable and $\mathcal{P}^{Z}$ is the Zwanzig projection operator as defined in eq. (28). Thus, the projection operator we chose in sec. IV.1 maximizes the drift term (i.e. the conservative term) and minimizes the rest (i.e. the dissipative term).
The crucial difference between sec. IV and ref. [14] is the definition of the inner product that induces the norm. The inner product employed in ref. [39] does not contain a weight, while the inner product required for the Hamiltonian of mean force to appear in eq. (48) contains the equilibrium measure $\varrho_{\beta}$. As demonstrated in ref. [44] a unique decomposition of a given generator can be achieved with respect to weighted scalar products. We thus agree with Colla and Breuer, one can use the generalized master equation and its unique splitting to define work. However, if the aim is a thermodynamic interpretation, we propose a different inner product, namely (29) with a canonical equilibrium distribution as weight. This yields the correct limit for systems coupled to a heat bath.